Project

General

Profile

NDOS CommissionMtng 20121210
Attending: Jon, Jan, Nick and Susan, Abbey (2 week), Sandeep, Leon

Jan - 14256 30 subrun (subrun zero 12 minute of data) (no empty events) (7153)(3 channel)
- Tuesday into Wednesday smooth running (subrun 9 is missing)
- 14359 Rough no empties sub run length issue (Raw -> Root issue)
- 14360 no empty
- 14363 (with flasher) empty events 40% empty events
- Sunday run 900 bad channels (2-2 60 is flaky)
- Header information (EVD info shows that run and subrun inconsisent)

Sandeep - Installed Flashers
- DCM 2-2 , DCM 2-3
- Still noisy

Leon - Test stand (taking noise data)
- Rebooting is needed because of a version conflict with DCMApp (FEB Firmware)
- RC Client crahsed once
- Leon tamed the bad apd.
- Continued test (shipping APD this week) 30 out tomorrow
- Leon will be here Thursday.
- Abbey and Sandeep are available.